Shift report 20000817 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 17 2000 - 08:01:28 EDT

  • Next message: Konstantin Olchanski: "Re: Shift report 20000817 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug 17 08:01:27 2000, Shift: 00:00-08:00
    Supervisor: HI, with: HI & CE
    Log book#4 pages 110 - 111
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from approx. 00:50. Intensity B=9416, Y=7390. 
       The beam was stable all night - Intensity at 08:00: 
       B~5800, Y~5700 (approx 10 events/sec).
    2. run002313   MRS at 90 deg. -4KGauss.
       run002314    "       "        "      (lost?)
       run002315    "       "        " 	
       Problem with the DAQ: At 04:20 we discovered that 
       the .dat file for a sequence was deleted from the 
       spool dir when a new sequence started. The DAQ 
       window gave no error messages. Only the .snd files 
       (it seems) are saved.
    3. T2a (L1458 1.6) anode HV tripped once around 03:30. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 17 2000 - 08:05:38 EDT